# Learning Efficient Representations of Neutrino Telescope Events

### v2

@@ Abstract
Neutrino telescopes detect rare interactions of particles produced in some of the most extreme environments in the Universe. This is accomplished by instrumenting a cubic-kilometer volume of naturally occurring transparent medium with light sensors. Given their substantial size and the high frequency of background interactions, these telescopes amass an enormous quantity of large variance, high-dimensional data. These attributes create substantial challenges for analyzing and reconstructing interactions, particularly when utilizing machine learning (ML) techniques. In this paper, we present a novel approach, called om2vec, that employs transformer-based variational autoencoders to efficiently represent neutrino telescope events by learning compact and descriptive latent representations. We demonstrate that these latent representations offer enhanced flexibility and improved computational efficiency, thereby facilitating downstream tasks in data analysis.
@@ 1 Introduction
Neutrino telescopes search for rare interactions caused by neutrinos, an elusive particle central to many open questions in the Standard Model of particle physics, such as the origin of their masses and the matter anti-matter asymmetry of the universe. The leading detector in operation today, the IceCube Neutrino Observatory, comprises 5,160 optical modules (OMs) arranged on strings deep within the clear ice of the Antarctic glacier (see [1]). Each OM is designed to precisely record the arrival times of photons, forming a photon arrival time distribution (PATD). The data gathered by IceCube are archived as "events," which are triggered when a specific number of coincident light detections occur between OMs in neighboring strings, signaling the presence of a charged particle. Figure 1 is an artistic representation of the photons produced in a neutrino interaction and their subsequent detection in the OMs. A neutrino telescope typically classifies events into categories based on the morphology of the light detected: cascade-like events, identifiable by their approximately spherical spatial distribution; and track-like events, characterized by their elongated, linear spatial morphology. In IceCube, events are recorded at an approximate rate of 3000 per second, constituting a large data rate. Moreover, comparable or even larger optical neutrino telescopes are either under construction or planned, such as KM3NeT [2], Baikal-GVD [3], P-ONE [4], IceCube-Gen2 [5], TRIDENT [6], and HUNT [7], signaling a need to develop methods for efficiently handling neutrino telescope data.
@@
The accurate reconstruction and prediction of the interacting particle’s physical properties—including its energy, direction, and type— from the PATDs recorded by each OM, is essential for informing and facilitating downstream physics analyses. A significant challenge in this task is the incorporation of the full-timing information provided by the PATDs of the OMs, which are characterized by high dimensionality and variable lengths. Interaction events may last several $\mu s$, yet ns-scale timing resolution is needed on the PATDs for most downstream physics analyses. Furthermore, as illustrated by OM#1 in Figure 1, some OMs near the interaction point can detect up to tens of thousands of photons in the highest-energy and brightest events. On the other hand, many further OMs record only a few photons, resulting in highly irregular events; illustrated in the OM#2 distribution on the same figure. In this paper, we introduce a transformer-based variational autoencoder (VAE), termed om2vec, which encodes PATDs into a fixed-size compact latent representation. om2vec is trained to reconstruct the input PATD as readout by OMs, using only the information from its encoded latent space. In this sense, it learns an efficient and more flexible representation of the data recorded in each OM. Our approach provides several significant advantages over existing methods, including improved information retention, greater reliability, faster processing speed, and reduced hyperparameter dependence. We also demonstrate that utilizing transformers for this particular problem is particularly effective, by comparing it to a baseline fully-connected network. We argue that these advantages make this the first “one-size-fits-all" neutrino event representation learner, making it suitable for deployment at the earliest stages of experimental data collection. This will serve to circumvent the challenges inherent in raw neutrino telescope data, and facilitate the rapid adoption of more sophisticated ML techniques from the image-processing community into the neutrino physics field. Source code, datasets, and pre-trained checkpoints can be found on GitHub.
@@ 1.1 Related Works
Machine learning techniques have successfully been applied to neutrino reconstruction tasks in recent years, notably in neutrino telescopes such as IceCube [8, 9, 10, 11, 12, 13, 14, 15] and KM3NeT [16, 17, 18, 19, 20]. Previous studies have employed various approaches to address the challenges of working with neutrino telescope data. One approach uses summary statistics [9], utilizing key variables that summarize the PATD of a given OM. This method yields greater efficiency and flexibility, as each OM has a fixed-size description of the timing distribution, allowing events to be formatted as 2D images with multiple features per pixel. However, the summary statistics lose a substantial amount of information present in the original PATD. An alternative method involves parameterizing the PATD by fitting an asymmetric Gaussian mixture model (AGMM). In this study, we implemented a basic AGMM to serve as a comparison, drawing inspiration from [10]. While this retains significantly more information than summary statistics, the optimization process can be slow and prone to failure, with a strong dependence on hyperparameters, e.g., the number of Gaussian components. Our transformer-based VAE approach builds upon these previous methods while addressing their limitations.
@@ 2 Architecture
Since their inception, VAEs have been extensively utilized for learning effective data representations and generative modeling [21]. More recently, transformers have demonstrated exceptional performance in handling sequential data [22]. Given the time series nature of our dataset, we incorporate transformer layers to enhance feature learning capabilities. Our model employs the standard encoder-decoder structure, as illustrated in Figure 2. We assume a Gaussian prior for the latent space and train the encoder to learn the parameters $\mu$ and $\sigma$, with some added random noise $\epsilon$. The re-parameterization trick is then utilized to construct the latent representation $z$ while maintaining proper gradient flow to these parameters. A key property of VAEs over regular autoencoders is their continuous latent space, meaning that similar representations within this space correspond to similar reconstructed PATDs.
@@ 3 Datasets
Simulations play a key role in neutrino physics experiments, facilitating preliminary testing that informs experiment development and subsequent physics analyses. As such, the field has developed sophisticated simulation tools capable of generating events for incorporation into our training datasets. In this paper, we present results trained and tested on simulated events, but emphasize that the methodology can be readily applied to real detector data. We employ the open-source simulation toolkit Prometheus [23] to emulate an IceCube-like array of OMs within Antarctic ice. We generate four datasets of events based on the different neutrino interaction types, corresponding to the distinct flavors of the neutrino, $\nu$. Specifically, these interactions are the charged-current electron neutrino ($\nu_{e}$ CC), muon neutrino ($\nu_{\mu}$ CC), and tau neutrino ($\nu_{\tau}$ CC) interactions, alongside the neutral-current ($\nu$ NC) interactions. Notably, the $\nu_{e}$ CC, $\nu_{\tau}$ CC, and $\nu$ NC interactions produce predominantly spherical cascade-like events, while the $\nu_{\mu}$ CC interactions result in long track-like events. Additionally, as further detailed in Section 5, the PATDs generated by $\nu_{\tau}$ CC interactions are of particular interest to physicists due to their potential to exhibit a smoking-gun “double-bang" signature of astrophysical neutrinos.
@@
Events are simulated, and all OMs detecting at least one photon hit are included in the overall training dataset. We extract five million PATDs from both $\nu_{e}$ CC and $\nu$ NC interactions, as their signals are quite similar. Additionally, 7.5 million PATDs are sourced from $\nu_{\mu}$ CC events, and 3.25 million PATDs from $\nu_{\tau}$ CC interactions, totaling approximately 20 million PATDs for the training dataset. An additional 4.5 million distributions are reserved for the test dataset, mixed roughly equally from each interaction type. The majority of OMs registering light detect only a few photon hits. For an illustration of this distribution, see the bottom panel in Figure 3. Notably, photon yields from interactions increase significantly at higher energies. The simulated events span a wide range of energies to encompass all types of PATDs that may occur in neutrino telescopes.
@@
As previously mentioned, each PATD in the datasets is pre-processed into a fixed-length vector by binning photon hits over time. A total of 6400 bins, each 1 ns wide, represent the first 6.4 $\mu s$ of hits—a readout time frame consistent with IceCube’s OMs [1]. The choice of binning and time window can be adapted to suit the specifications of other neutrino telescopes. Any remaining hits are accumulated in a final overflow bin. The hit counts in each bin are then log-normalized, as they can be extremely large for high-energy neutrino interactions.
@@ 4 Proposed Methods
[…]
Given the predicted PDF from the network (after softmax), and the normalized true PDF from the input PATD, this loss function takes the negative logarithm of the summed Poisson likelihoods, representing the probability of registering the observed number of hits in each bin, based on the predicted hit counts.
The total training loss combines reconstruction loss and a KL divergence regularization loss on the latent space. This regularization term is needed to ensure that the learned latent space aligns with the prior normal distribution. It is also scaled by a $\beta$ factor that follows a cyclic cosine function during training, peaking at a hyperparameter value set at $10^{-5}$. A batch size of 1024 is used during training.
@@ 5 Results
$AG(t\mid\mu,\sigma,r)=\frac{2}{\sqrt{2\pi}*\sigma(r+1)}*\begin{cases}\exp({-\frac{(t-\mu)^{2}}{2\sigma^{2}}}),&x\leq\mu\\ \exp({-\frac{(t-\mu)^{2}}{2(r\sigma)^{2}}}),&\text{otherwise}\end{cases}$
where $\mu$ and $\sigma$ are the mean and standard deviation and $r$ is the asymmetry parameter, as introduced in [10]. Each asymmetric Gaussian is assigned a weighting factor, $w$, and then summed. Consequently, each component of the mixture model introduces four parameters: $\mu$, $\sigma$, $r$, and $w$. The mixture model is initially fit using $k$-means clustering to obtain the starting parameters, followed by solving a bounded-constrained optimization problem that minimizes the negative log-likelihood. However, the fit is only performed when the number of photon hits (the length of $x$) is greater than the number of Gaussian components. Otherwise, the true photon hit times are stored.
@@ 5.1 Timing Distribution Reproducibility
Figure 3: JS distance between the true input PATD and the reconstructed PATD across different methods, plotted as a function of the number of detected photons in the input PATD. Greater values of the JS distance implies a worse fit. In the AGMM methods, the sudden jump is caused by the fit only being performed when the number of photon hits is greater than the number of Gaussian components. The lower panel illustrates the percentage of PATDs in each number of photons bin, relative to the total dataset.
@@
To assess the reconstruction capability, we employ the Jensen-Shannon (JS) distance [24, 25], which is derived from the KL divergence and is defined as follows:
@@
given the distributions $P$ and $Q$, and where $D_{KL}$ is the KL divergence. The JS distance provides a similarity score between 0 and 1, where 0 indicates identical distributions and 1 indicates maximum difference. In Figure 3, we present the median JS distance curves for both om2vec and AGMM approaches, across the three different latent dimension sizes. We plot the JS distance as a function of the number of photons recorded in the PATD, scaling up to a thousand photons. While some PATDs observe significantly more photons from the highest-energy events, the statistics are sparse (i.e., there are few PATDs) at those higher photon counts.
@@
Figure 4: Failure percentage ($>0.99$ JS distance) as a function of the number of photons on the input PATD. om2vec is always at 0% in this figure.
@@
Notable characteristics of these curves include well-reconstructed distributions on the left, where the number of photons is low enough for the representations to “memorize” the distribution. This is followed by a sharp upward trend, peaking in a regime where there are too many photons for memorization, but insufficient statistics to generate a well-formed PATD. We then observe the JS distance decreasing as the number of photons increases, allowing for enough statistical information to effectively reconstruct the PATD. We also observe that increasing the latent dimension size for om2vec noticeably enhances the performance of representation learning, a trend not reflected in the AGMM method. Further insights into this phenomenon can be gained by examining the failure rate in Figure 4, where we define “failures” as PATDs reconstructed with a JS distance greater than 0.99. Notably, the VAE is never above this threshold for all PATDs. In contrast, a substantial percentage of PATDs from the AGMM method are mis-reconstructed due to optimization failures, with this failure rate increasing as more parameters and components are added.
@@
Figure 5: An example PATD of a “double-bang" signature, characterized by two distinct peaks. om2vec is able to reconstruct both peaks, while the AGMM is noticeably less accurate.
@@
In order to give a more intuitive sense of the performance of om2vec, we examine individual PATDs and their reconstructed probability density functions. This is especially interesting for exotic cases such as the “double-bang” signature produced by $\nu_{\tau}$ interactions, as mentioned earlier. Figure 5 illustrates an example of this type of PATD, distinguished by its two distinct peaks. It is evident that om2vec reconstructs the bimodal structure of the timing distribution with greater precision than the AGMM method, despite the overall JS distance difference between the two methods being minor. As this is an extremely large PATD ($\sim$100,000 photons), this is likely due to the dominant statistics contained in the first peak. However, from a physics perspective, retaining information about both peaks is crucial for downstream tasks that attempt to isolate $\nu_{\tau}$ interactions, a smoking-gun indication of astrophysical neutrinos.
@@
Table 1: Average JS distance and forward-pass FLOPs for the different om2vec models tested.
@@
We further evaluated three different model architectures, all employing 64 latent parameters, to assess their impact on the average JS distance. As a baseline case, we trained a model using only fully-connected layers. The other two architectures incorporated transformers, as it is described in Section 2, with varying numbers of encoder and decoder blocks. A summary of the results is provided in Table 1. We note that the default architecture used in previous experiments features three encoder-decoder blocks and is highlighted in bold. Comparing the fully connected model to the transformer-based architectures, we observe a substantial improvement in reconstruction performance when incorporating transformer layers. However, this enhancement comes at a considerable computational cost, with the number of FLOPs required for a forward pass increasing by approximately two orders of magnitude. Furthermore, adding additional transformer encoder-decoder layers yields a slight further improvement in reconstruction performance.
@@
Figure 6: Angular resolution as a function of the true neutrino energy. The models trained using the latent representations from om2vec are able to perform just as well when compared to using the full timing information.
We now evaluate whether downstream physics analysis tasks suffer any performance loss when using the latent representations. One of the most critical downstream analysis tasks is angular reconstruction, i.e. predicting the direction of the incoming particle based on the photons recorded across the OMs. We test three different separately trained ML models for angular reconstruction on track-like events.
  • SSCNN (Full): a sparse submanifold convolutional neural network (SSCNN) that utilizes full timing information in a 4D CNN, as described in [12].
@@
In the methods employing om2vec, we first pre-process the events using the trained 64-parameter om2vec model, to generate latent representations for the subsequent SSCNN and CNN networks. In Figure 6, we show the neutrino angular resolution (the angular difference between the true and predicted neutrino direction), as a function of the true neutrino energy. We find that the SSCNN models using the full timing information and the latent representations perform similarly, though as expected, the full 4D model is able to achieve a slightly better resolution. Notably, SSCNN sacrifices some performance to remain computationally feasible on higher-dimensional sparse data [26]. Thus, we observe that the dense CNN approach with latent representations performs just as well or slightly better than the SSCNN with full-timing information, depending on the true neutrino energy. This CNN approach is made possible by the fixed-length latent representations. Importantly, we note that this new flexibility allows for the easy adaptation of more powerful models, such as vision transformers [27], for angular reconstruction.

### v3

@@ Abstract
Neutrino telescopes detect rare interactions of particles produced in some of the most extreme environments in the Universe. This is accomplished by instrumenting a cubic-kilometer scale volume of naturally occurring transparent medium with light sensors. Given their substantial size and the high frequency of background interactions, these telescopes amass an enormous quantity of large variance, high-dimensional data. These attributes create substantial challenges for analyzing and reconstructing interactions, particularly when utilizing machine learning (ML) techniques. In this paper, we present a novel approach, called om2vec, that employs transformer-based variational autoencoders to efficiently represent the detected photon arrival time distributions of neutrino telescope events by learning compact and descriptive latent representations. We demonstrate that these latent representations offer enhanced flexibility and improved computational efficiency, thereby facilitating downstream tasks in data analysis.
@@ 1 Introduction
Neutrino telescopes search for rare interactions caused by neutrinos, elusive particles central to many open questions in the Standard Model of particle physics, such as the origin of their masses and the matter anti-matter asymmetry of the universe. The leading detector in operation today, the IceCube Neutrino Observatory, comprises 5,160 optical modules (OMs) arranged on strings deep within the clear ice of the Antarctic glacier (see [IceCube:2016zyt]). Each OM is designed to precisely record the arrival times of photons, forming a photon arrival time distribution (PATD). The data gathered by IceCube are archived as "events," which are triggered when a specific number of coincident light detections occur between OMs in neighboring strings, signaling the presence of a charged particle. Figure˜1 is an artistic representation of the photons produced in a neutrino interaction and their subsequent detection in the OMs. A neutrino telescope typically classifies events into categories based on the morphology of the light detected: cascade-like events, identifiable by their approximately spherical spatial distribution; and track-like events, characterized by their elongated, linear spatial morphology. In IceCube, events are recorded at an approximate rate of 3000 per second, constituting a large data rate. Moreover, comparable or even larger optical neutrino telescopes are either under construction or planned, such as KM3NeT [KM3NeT:2018wnd], Baikal-GVD [Baikal-GVD:2018isr], P-ONE [P-ONE:2020ljt], IceCube-Gen2 [IceCube-Gen2:2020qha], TRIDENT [Ye:2022vbk], and HUNT [Huang:2023mzt], signaling a need to develop methods for efficiently handling neutrino telescope data.
@@
The accurate reconstruction and prediction of the interacting particle’s physical properties—including its energy, direction, and type— from the PATDs recorded by each OM, is essential for informing and facilitating downstream physics analyses. A significant challenge in this task is the incorporation of the full-timing information provided by the PATDs of the OMs, which are characterized by high dimensionality and variable lengths. Interaction events may last several $\mu s$, yet ns-scale timing resolution is needed on the PATDs for most downstream physics analyses. Furthermore, as illustrated by OM#1 in Figure˜1, some OMs near the interaction point can detect up to tens of thousands of photons in the highest-energy and brightest events. On the other hand, many further OMs record only a few photons, resulting in highly irregular events; illustrated in the OM#2 distribution on the same figure. In this paper, we introduce a transformer-based variational autoencoder (VAE), termed om2vec, which encodes PATDs into a fixed-size compact latent representation. om2vec is trained to reconstruct the input PATD as readout by OMs, using only the information from its encoded latent space. In this sense, it learns an efficient and more flexible representation of the data recorded in each OM. Our approach provides several significant advantages over existing methods, including improved information retention, greater reliability, faster processing speed, and reduced hyperparameter dependence. We also demonstrate that utilizing transformers for this particular problem is particularly effective, by comparing it to a baseline fully-connected network. We argue that these advantages make this the first “one-size-fits-all" neutrino event representation learner, making it suitable for deployment at the earliest stages of experimental data collection. This will serve to circumvent the challenges inherent in raw neutrino telescope data, and facilitate the rapid adoption of more sophisticated ML techniques from the image-processing community into the neutrino physics field. Source code, datasets, and pre-trained checkpoints can be found on GitHub.
@@ 1.1 Related Works
Machine learning techniques have successfully been applied to neutrino reconstruction tasks in recent years, notably in neutrino telescopes such as IceCube [IceCube:2018gms, Abbasi:2021ryj, IceCube:2021umt, IceCube:2022njh, PhysRevD.108.063017, Eller:2023myr, Jin:2023xts, Zhu:2024ubz] and KM3NeT [DeSio:2019lcr, KM3NeT:2020zod, Reck:2021zqw, Guderian:2022cco, Mauro:2023zam]. Previous studies have employed various approaches to address the challenges of working with neutrino telescope data. One approach uses summary statistics [Abbasi:2021ryj], utilizing key variables that summarize the PATD of a given OM. This method yields greater efficiency and flexibility, as each OM has a fixed-size description of the timing distribution, allowing events to be formatted as 2D images with multiple features per pixel. However, the summary statistics lose a substantial amount of information present in the original PATD. An alternative method involves parameterizing the PATD by fitting an asymmetric Gaussian mixture model (AGMM). In this study, we implemented a basic AGMM to serve as a comparison, drawing inspiration from [IceCube:2021umt]. While this retains significantly more information than summary statistics, the optimization process can be slow and prone to failure, with a strong dependence on hyperparameters, e.g., the number of Gaussian components. Our transformer-based VAE approach builds upon these previous methods while addressing their limitations.
@@ 2 Architecture
Since their inception, VAEs have been extensively utilized for learning effective data representations and generative modeling [kingma2022autoencodingvariationalbayes]. More recently, transformers have demonstrated exceptional performance in handling sequential data [NIPS2017-3f5ee243]. Given the time series nature of our dataset, we incorporate transformer layers to enhance feature learning capabilities. Our model employs the standard encoder-decoder structure, as illustrated in Figure˜2. We assume a Gaussian prior for the latent space, and train the encoder to predict the parameters $\mu$ and $\sigma$, which correspond to the mean and standard deviation, respectively, of the approximate posterior distribution over the latent variables. These parameters describe the distribution from which the latent representation is sampled, incorporating randomness via a noise variable $\epsilon$. The re-parameterization trick [kingma2013auto] is then utilized to construct the latent representation $z=\mu+\sigma\cdot\epsilon$, allowing for proper gradient flow to $\mu$ and $\sigma$ during training. A key property of VAEs over regular autoencoders is their continuous latent space, meaning that similar representations within this space correspond to similar reconstructed PATDs.
@@ 3 Datasets
Simulations play a key role in neutrino physics experiments, facilitating preliminary testing that informs experiment development and subsequent physics analyses. As such, the field has developed sophisticated simulation tools capable of generating events for incorporation into our training datasets. In this paper, we present results trained and tested on simulated events, but emphasize that the methodology can be readily applied to real detector data. We employ the open-source simulation toolkit Prometheus [LAZAR2024109298] to emulate an IceCube-like array of OMs within Antarctic ice. Prometheus enables the simulation of neutrino telescope events from particle injection to detector response across arbitrary detector media and geometries. However, it does not incorporate experiment-specific detector effects that are not publicly available. We generate four datasets of events based on the different neutrino interaction types, corresponding to the distinct flavors of the neutrino, $\nu$. Specifically, these interactions are the charged-current electron neutrino ($\nu_{e}$ CC), muon neutrino ($\nu_{\mu}$ CC), and tau neutrino ($\nu_{\tau}$ CC) interactions, alongside the neutral-current ($\nu$ NC) interactions. Notably, the $\nu_{e}$ $C$, $\nu_{\tau}$ CC, and $\nu$ NC interactions produce predominantly spherical cascade-like events, while the $\nu_{\mu}$ CC interactions result in long track-like events. Additionally, as further detailed in Section˜5, the PATDs generated by $\nu_{\tau}$ CC interactions are of particular interest to physicists due to their potential to exhibit a smoking-gun “double-bang" signature of astrophysical neutrinos.
@@
Events are simulated, and all OMs detecting at least one photon hit are included in the overall training dataset. We extract five million PATDs from both $\nu_{e}$ CC and $\nu$ NC interactions, as their signals are quite similar. Additionally, 7.5 million PATDs are sourced from $\nu_{\mu}$ CC events, and 3.25 million PATDs from $\nu_{\tau}$ CC interactions, totaling approximately 20 million PATDs for the training dataset. An additional 4.5 million distributions are reserved for the test dataset, mixed roughly equally from each interaction type. The majority of OMs registering light detect only a few photon hits. For an illustration of this distribution, see the bottom panel in Figure˜3. Notably, photon yields from interactions increase significantly at higher energies. The simulated events span a wide range of energies from $100$ GeV to $1$ PeV to encompass all types of PATDs that may occur in neutrino telescopes.
@@
As previously mentioned, each PATD in the datasets is pre-processed into a fixed-length vector by binning photon hits over time. A total of 6400 bins, each 1 ns wide, represent the first 6.4 $\mu s$ of hits—a readout time frame consistent with IceCube’s OMs [IceCube:2016zyt]. The choice of binning and time window can be adapted to suit the specifications of other neutrino telescopes. Any remaining hits are accumulated in a final overflow bin. The hit counts in each bin are then log-normalized, as they can be extremely large ($\sim$10,000) for high-energy neutrino interactions.
@@ 4 Proposed Methods
$n_{i}$ $=\text{observed count in bin }i\text{ (true PDF)}$
$\lambda_{i}(\theta)$ $=\text{expected count in bin }i\text{ (predicted PDF)}$
$N_{\text{bins}}$ $=\text{total number of bins, or input length}$
$\theta$ $=\text{model parameters}$
Given the predicted PDF from the network (after the use of softmax), and the normalized true PDF from the input PATD, this loss function takes the negative logarithm of the summed Poisson likelihoods, representing the probability of registering the observed number of hits in each bin, based on the predicted hit counts.
The total training loss combines the reconstruction loss and a KL divergence regularization loss on the latent space. This regularization term is needed to ensure that the learned latent space aligns with the prior normal distribution. It is also scaled by a $\beta$ factor that follows a cyclic cosine function during training, peaking at a hyperparameter value set at $10^{-5}$. A batch size of 1024 is used during training.
@@ 5 Results
$AG(t\mid\mu,\sigma,r)=\frac{2}{\sqrt{2\pi}*\sigma(r+1)}*\begin{cases}\exp({-\frac{(t-\mu)^{2}}{2\sigma^{2}}}),&x\leq\mu\\ \exp({-\frac{(t-\mu)^{2}}{2(r\sigma)^{2}}}),&\text{otherwise}\end{cases}$ (5.1)
where $\mu$ and $\sigma$ are the mean and standard deviation and $r$ is the asymmetry parameter, as introduced in [IceCube:2021umt]. Each asymmetric Gaussian is assigned a weighting factor, $w$, and then summed. Consequently, each component of the mixture model introduces four parameters: $\mu$, $\sigma$, $r$, and $w$. The mixture model is initially fit using $k$-means clustering to obtain the starting parameters, followed by solving a bounded-constrained optimization problem that minimizes the negative log-likelihood. However, the fit is only performed when the number of photon hits (the length of $x$) is greater than the number of Gaussian components. Otherwise, the true photon hit times are stored.
@@ 5.1 Timing Distribution Reproducibility
Figure 3: Jensen-Shannon distance between the true input PATD and the reconstructed PATD across different methods, plotted as a function of the number of detected photons in the input PATD. Greater values of the JS distance implies a worse fit. In the AGMM methods, the sudden jump is caused by the fit only being performed when the number of photon hits is greater than the number of Gaussian components. The lower panel illustrates the percentage of PATDs in each number of photons bin, relative to the total dataset.
@@
To assess the reconstruction capability, we employ the Jensen-Shannon (JS) distance [61115, e21050485], which is derived from the KL divergence and is defined as follows:
@@
given the distributions $P$ and $Q$, and where $D_{KL}$ is the KL divergence. The JS distance provides a similarity score between 0 and 1, where 0 indicates identical distributions and 1 indicates maximum difference. In Figure˜3, we present the median JS distance curves for both om2vec and AGMM approaches, across the three different latent dimension sizes. We plot the JS distance as a function of the number of photons recorded in the PATD, scaling up to a thousand photons. While some PATDs observe significantly more photons from the highest-energy events, the statistics are sparse (i.e., there are few PATDs) at those higher photon counts.
@@
Figure 4: Failure percentage (>0.99>0.99 JS distance) as a function of the number of photons on the input PATD. om2vec is always at 0% in this figure.
@@
Notable characteristics of these curves include well-reconstructed distributions in the first few bins, where the number of photons is low enough for the representations to “memorize” the distribution. This is followed by a sharp upward trend, culminating in a regime where the photon count exceeds the memorization capacity (typically around 30–70 hits for the om2vec models), yet remains too low to yield a well-formed PATD due to insufficient statistics. We then observe the JS distance decreasing as the number of photons increases, allowing for enough statistical information to effectively reconstruct the PATD. We also observe that increasing the latent dimension size for om2vec noticeably enhances the performance of representation learning, a trend not reflected in the AGMM method. Further insights into this phenomenon can be gained by examining the failure rate in Figure˜4, where we define “failures” as PATDs reconstructed with a JS distance greater than 0.99. Notably, the VAE is never above this threshold for all PATDs. In contrast, a substantial percentage of PATDs from the AGMM method are mis-reconstructed due to optimization failures, with this failure rate increasing as more parameters and components are added. Since the VAE framework is self-supervised, it can be trained directly on real detector data, suggesting that with sufficient statistics, it could adapt to simulation–data discrepancies and reliably reconstruct PATDs even on data.
@@
Figure 5: Two example PATDs, the bottom one being a “double-bang" signature, characterized by two distinct peaks. om2vec is able to reconstruct both peaks, while the AGMM is noticeably less accurate.
@@
In order to give a more intuitive sense of the performance of om2vec, we examine individual PATDs and their reconstructed probability density functions. This is especially interesting for exotic cases such as the “double-bang” signature produced by $\nu_{\tau}$ interactions, as mentioned earlier. Figure˜5 illustrates an example of this type of PATD, distinguished by its two distinct peaks. It is evident that om2vec reconstructs the bimodal structure of the timing distribution with greater precision than the AGMM method, despite the overall JS distance difference between the two methods being minor. As this is an extremely large PATD ($\sim$100,000 photons), this is likely due to the dominant statistics contained in the first peak. However, from a physics perspective, retaining information about both peaks is crucial for downstream tasks that attempt to isolate $\nu_{\tau}$ interactions, a smoking-gun indication of astrophysical neutrinos.
We further evaluated three different model architectures, all employing 64 latent parameters, to assess their impact on the average JS distance. As a baseline case, we trained a model using only fully-connected layers. The other two architectures incorporated transformers, as it is described in Section˜2, with varying numbers of encoder and decoder blocks. A summary of the results is provided in Table˜1. We note that the default architecture used in previous experiments features three encoder-decoder blocks and is highlighted in bold. By comparing the fully connected model to the transformer-based architectures, we observe a substantial improvement in reconstruction performance when incorporating transformer layers. However, this enhancement comes with a computational cost, as the number of floating-point operations per second (FLOPs), which quantify computational expense, increases by approximately two orders of magnitude during a forward pass. Furthermore, adding additional transformer encoder-decoder layers yields a slight further improvement in reconstruction performance.
@@
Table 1: Average JS distance (JSD) and forward-pass FLOPs for the different om2vec models tested.
@@
Figure 6: Angular resolution as a function of the true neutrino energy. The shaded regions show the $20^{\text{th}}$ to $80^{\text{th}}$ percentile bands. The models trained using the latent representations from om2vec are able to perform just as well when compared to using the full timing information.
We now evaluate whether downstream physics analysis tasks suffer any performance loss when using the latent representations, and whether these representations offer additional benefits such as faster reconstruction runtime, as discussed in Section˜5.3. One of the most critical downstream analysis tasks is angular reconstruction, i.e. predicting the direction of the incoming particle based on the photons recorded across the OMs. We test three different separately trained ML models for angular reconstruction on track-like events.
  • SSCNN (Full): a sparse submanifold convolutional neural network (SSCNN) that utilizes full timing information in a 4D CNN, as described in [PhysRevD.108.063017].
@@
In the methods employing om2vec, we first pre-process the events using the trained 64-parameter om2vec model, to generate latent representations for the subsequent SSCNN and CNN networks. In Figure˜6, we show the neutrino angular resolution (the angular difference between the true and predicted neutrino direction), as a function of the true neutrino energy. We find that the SSCNN models using the full timing information and the latent representations perform similarly, though as expected, the full 4D model is able to achieve a slightly better resolution. Notably, SSCNN sacrifices some performance to remain computationally feasible on higher-dimensional sparse data [sscnn]. Thus, we observe that the dense CNN approach with latent representations performs just as well or slightly better than the SSCNN with full-timing information, depending on the true neutrino energy. This CNN approach is made possible by the fixed-length latent representations, which also reduce the input data dimensionality and simplify the model pipeline. Importantly, we note that this new flexibility allows for the easy adaptation of more powerful models, such as vision transformers [dosovitskiy2021an], for angular reconstruction.